If the length of a rectangle is 8 more than it's width, what is the area of the rectangle?
Answer by Theo(13342)

area of a rectangle is equal to L * W
If the Length is 8 more than the width, then:
L = W + 8 and the area of the rectangle is equal to (W + 8) * W.
The area of the rectangle becomes W^2 + 8W.
There is any number of values that will satisfy this equation.
Just pick a value for W and you can find the area of the rectangle and you can find the length of the rectangle as well.
for example:
let W = 5.
The area of the rectangle is 25 + 40 = 65 square units.
The length of the rectangle is equal to 5 + 8 = 13.
L * W = 5 * 13 = 65 square units.
Let W equal to 1 and the area of the rectangle is equal to 1^2 + 8*1 = 9 square units.
the width is 1 and the length is 1 + 8 = 9.
L * W = 1 * 9 = 9 square units of area.
You do not have enough information to find only 1 area of the rectangle.
If you were told that the length or the width was a certain number of units, then you could find the 1 area that would solve the problem.
As it stands now, your area, and your length, are defined by the value of W that you choose.